Please answer part A,B,and C. Also, please explain your steps as you work through the problem. thanks.
(A) Caculat te reccion Corux AN 25hw.
Answer #1

3 kN/m 3. YM eam ABC 15kN 3 KN/m 6 2.2S m Cy(+) + 25 =15x2 Beam ASC .. 3(4-5+1-5) +1.25 Ayt By = By- 13-667 kN () Ay 5.583 kNca* at Point E Intei NE 2-25 m Ay5 S83 kN 2-2S ME= S-583 (2.25 ME = 4.368 kN-m 5 kN Intennal foarcA at PontF F C 2m , Dyー13.T